OpenStudy (dtan5457):

Find the other 5 trig functions if cos=5/6 domain=3pi/2

OpenStudy (dtan5457):

Someone refresh me on this again..

OpenStudy (jdoe0001):

ok so based on that domain wer're on the 4th quadrant meaning, cosine is positive, sine is negative, tangent is .. .negative becasue the other two differ in signs anyway hypotenuse is radius, is always positive

OpenStudy (dtan5457):

Ohhh. I should solve for sin using a identity..

OpenStudy (dtan5457):

For a second I thought I was suppose to use the unit circle somehow..

OpenStudy (dtan5457):

sin^2+(5/6)^2=1 sin^2=sqrt11/6

OpenStudy (dtan5457):

well, not sin^2 but sin

OpenStudy (dtan5457):

then, i just use cos/sin for tangent, and inverse everything right?

OpenStudy (dtan5457):

then I just make sure only cos/sec are positive?

OpenStudy (jdoe0001):

hmm actually lemme fix that too =)

OpenStudy (jdoe0001):

hmmm wait a sec.. man l hold

OpenStudy (dtan5457):

I don't know, I remember solving these with identities. Maybe you can check my work? cos=5/6 sin=-sqrt11/6 tan=-5(sqrt11)/11 sec=6/5 csc=-6(sqrt11)/11 cot=-sqrt11/5

OpenStudy (jdoe0001):

\(\bf {\color{blue}{ cos}}(\theta)=\cfrac{adjacent}{hypotenuse} \to \cfrac{5}{6}\to \cfrac{a=5}{c=6} \\ \quad \\ c^2=a^2+{\color{blue}{ b}}^2\implies \pm\sqrt{c^2-a^2}={\color{blue}{ b}}\impliedby \textit{we know sine is negative} \\ \quad \\ -\sqrt{c^2-a^2}={\color{blue}{ b}}\qquad sin(\theta)=\cfrac{{\color{blue}{ b}}}{c}\qquad tan(\theta)=\cfrac{{\color{blue}{ b}}}{a}\)

OpenStudy (jdoe0001):

well... yes.. just that tangent is b/a a = 5

OpenStudy (dtan5457):

I got tangent wrong?

OpenStudy (jdoe0001):

well... tangent \(\bf tan(\theta)=\cfrac{opposite}{adjacent}\)

OpenStudy (dtan5457):

I just put cos/sin, 5/6 times 6/radical 11

OpenStudy (jdoe0001):

adjancent = 5 hypotenuse = 6 opposite = \(-\sqrt{11}\)

OpenStudy (dtan5457):

does that mean my sin and cos are also wrong

OpenStudy (jdoe0001):

nope, your cosine and sine are correct, but you have -> tan=-5(sqrt11)/11 <--

OpenStudy (jdoe0001):

\(\bf {\color{brown}{ sin}}(\theta)=\cfrac{opposite}{hypotenuse} \qquad \qquad % cosine {\color{blue}{ cos}}(\theta)=\cfrac{adjacent}{hypotenuse} \\ \quad \\ % tangent tan(\theta)=\cfrac{opposite}{adjacent} \qquad \qquad % cotangent cot(\theta)=\cfrac{adjacent}{opposite} \\ \quad \\ % cosecant csc(\theta)=\cfrac{hypotenuse}{opposite} \qquad \qquad % secant sec(\theta)=\cfrac{hypotenuse}{adjacent}\)

OpenStudy (dtan5457):

\[\frac{ 5 }{ 6 }\times -\frac{ 6 }{ \sqrt{11} }=\frac{ -30 }{ 6\sqrt{11} }=\frac{ -5 }{ \sqrt{11} }=\frac{ -5\sqrt{11} }{ 11 }\] let's say we just used tan=cos/sin, what did I do wrong?

OpenStudy (jdoe0001):

\(\bf tan(\theta)=\cfrac{opposite}{adjacent}\to \cfrac{\frac{-\sqrt{11}}{6}}{\frac{5}{6}}\to \cfrac{-\sqrt{11}}{\cancel{6}}\cdot \cfrac{\cancel{6}}{5}\)

OpenStudy (dtan5457):

lol you should of just told me tan=sin/cos, not cos/sin xD

OpenStudy (jdoe0001):

hmmm well.... yes tan = sin /cos :) " cos=5/6 <----- sin=-sqrt11/6 <----- tan=-5(sqrt11)/11 sec=6/5 csc=-6(sqrt11)/11 cot=-sqrt11/5 "

OpenStudy (dtan5457):

yeah i see that now. tan=-sqrt11/5 cot=-5sqrt11/11 basically a switch

OpenStudy (jdoe0001):

yeap

OpenStudy (jdoe0001):

cot and csc and sec are just the other two upside-down

OpenStudy (jdoe0001):

the other three rather

OpenStudy (dtan5457):

thank you :)

OpenStudy (jdoe0001):

yw